# Foliations by Critical Surfaces of the Hawking Energy in Asymptotically Flat Initial Data Sets

### v1

@@ 3. Center of the foliation
(71) $\pi(\tilde{\rho},\tilde{\rho})=o(|\tilde{x}|^{-2}),\quad(\nabla_{\tilde{\rho}}\pi)(\tilde{\rho},\tilde{\rho})=o(|\tilde{x}|^{-3}),\quad k(\cdot,\tilde{\rho})=o(|\tilde{x}|^{-2})$
and
(72) $|k^{\text{even}}|(\tilde{x})+|\tilde{x}||(\nabla k)^{\text{odd}}|(\tilde{x})=\mathcal{O}(|\tilde{x}|^{-3})$
@@
(76) $\begin{split}|\xi|^{-1}\xi^{i}(\partial_{i}G_{r,3})(\xi)=&-r^{2}\int_{S_{r}(q)}2W_{2}\alpha d\mu_{\delta}-r^{2}\int_{S_{r}(q)}2DW_{2}\left(r\delta(\frac{\xi}{|\xi|},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}\\ &-\frac{r^{2}}{2}\bar{D}^{2}\left(\int_{S_{r}(\xi r)}2W_{2}\alpha d\mu_{\delta}\right)_{|\xi=\tilde{\xi}}(r\xi-q,r\xi-q).\\ =&\mathcal{O}(r^{-1})+o(r^{-1}).\end{split}$
@@
Now proceeding as before and using that $\frac{r^{3}}{128\pi}\int_{S_{r}(q)}2DW_{2}\left(r\delta(\frac{\xi}{|\xi|},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}=o(1)$ we have
(77) $\begin{split}r\delta(\xi(r),&e_{i})=\delta(C_{H},e_{i})+\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\mathrm{Sc}\,\delta(e_{i},\nu)+r\pi(\nu,\nu)\nabla_{\nu}\pi(\nu,\nu)\delta(\nu,e_{i})+2\pi(\nu,\nu)k(\nu,e_{i})\\ &+\pi(\nu,\nu)^{2}\delta(\nu,e_{i})\,d\mu_{\delta}+\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\Big{(}\delta(e_{i},\nu)\delta(r\xi-q,\nu)\bar{D}_{\nu}\mathrm{Sc}+3\delta(e_{i},\nu)\delta(r\xi-q,\nu)\frac{\mathrm{Sc}}{r}\\ &-\frac{\mathrm{Sc}}{r}\,\delta(e_{i},r\xi-q)\Big{)}d\mu_{\delta}+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}2DW_{2}\left(\delta(e_{i},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}+\mathcal{O}(r^{-1})\\ &\,{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}11}=\delta(C_{H}+Z,e_{i})+\mathcal{O}(r^{-1})+o(1)\end{split}$
@@
and
(80) $|k^{\text{even}}|(\tilde{x})+|\tilde{x}||(\nabla k)^{\text{odd}}|(\tilde{x})=o(|\tilde{x}|^{-3})$
@@
(81) $\lim_{r\to\infty}\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\mathrm{Sc}\,\nu^{j}+2rDW_{2}\left(\delta(e_{j},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}=Z^{j}$
@@ 4.1. Nonnegativity along the foliation
In [27], the following nonnegativity and rigidity result for Hawking surfaces was obtained.
Theorem 4.1 ([27, Theorem 3.6]).
@@
As mentioned in the remark following [27, Theorem 3.6] one could define $f$ differently,
@@
First, we note that the Hawking energy is positive both along the foliation of Theorem 2.10 and for the on–center family of Theorem 2.14. This follows directly from [26, Theorem 2.2] (which treats the foliation) or from a direct application of Theorem 4.1.
@@ 4.2. Monotonicity
Monotonicity lies at the very heart of what one expects from a physically meaningful quasi-local energy: as a family of surfaces expands outward in an initial data set satisfying the dominant energy condition, its measured energy should not decrease. One of the most celebrated applications of monotonicity properties is in the proof of the Riemannian Penrose inequality by Huisken and Ilmanen [16]. In their work, they employed the inverse mean curvature flow to show that the Hawking energy is nondecreasing along the flow under suitable conditions, such as nonnegative scalar curvature. This monotonicity, in the time-symmetric initial data setting, allowed them to rigorously relate the ADM mass of an asymptotically flat manifold to the area of its outermost apparent horizon.
Since then, there have been several generalizations of this monotonicity for the Hawking energy to different settings. For instance, Bray, Jauregui, and Mars in [2, 3] established monotonicity results for the spacetime setting for "time flat" surfaces, while more recently, Hirsch studied monotonicity in a general initial data set setting [14].
@@
We proceed as in [20, Theorem 4]. We compute the variation of the Hawking energy
@@
In the early development of the conformal method for constructing solutions to the Einstein constraint equations, York studied the asymptotically flat case in [32] and proposed the following model for the extrinsic curvature:
@@
When deriving this model in Chapter 6 of York’s paper [32], the transverse-traceless (TT) part of the tensor $A$, denoted by $A^{*}$, is not explicitly determined during the construction of the model for asymptotically flat manifolds. Instead, York focuses on the decomposition of $A$ into its TT part and a divergence-free component derived from the vector potential $W$:
@@ 4.3. Rigidity on the foliation
In [27], the following rigidity result for the Wilmore foliation was obtained.
Theorem 4.11 ([27, Theorem 2.16]).
@@
Denoting by $\Omega$ interior of $\Sigma$, we can directly apply Theorem 4.1 obtaining that $\Omega$ is isomeric to a spacelike hypersurface in Minkowski spacetime, also we have that $\Sigma$ is a round sphere with $|k|_{\Sigma}=\lambda=|\mathring{B}|=\mathrm{Sc}_{|\Sigma}=\mathrm{Ric}_{|\Sigma}(\nu,\nu)=0$ and $H=\frac{2}{r^{2}}$ is constant, where $r$ is the area radius of $\Sigma$ (for more details see [27, Theorem 3.6]). Then what it remains to see is that $M\setminus\Omega$ is also isometric to a hypersurface in Minkowski space.
@@ Appendix A Density of Harmonic asymptotics
Corvino and Schoen showed in [6, Theorem 1] that initial data sets with harmonic asymptotics are dense in the set of solutions of the set of asymptotically flat vacuum solutions of the constrained equations. This result was later generalized to solutions of the Einstein constraint equations under the dominant energy condition by Eichmair, Huang and Lee in [7, Theorem 18], this result can be stated assuming higher regularity like in Definition A.3 (the proof of the result is basically unchanged).
@@
Note that in particular, we have also a density result in the Hölder norm by the following result which can be found in [21, Theorem A.25]:
@@ Acknowledgements.
The Author would like to thank Jan Metzger, Marc Mars, Rodrigo Avalos and Thomas Koerber for the helpful comments and discussions about this work. This research received partial support from the DFG as part of the SPP 2026 Geometry at infinity, project ME 3816/3-1.

### v2

@@ 3. Center of the foliation
(83) $\pi(\tilde{\rho},\tilde{\rho})=o(|\tilde{x}|^{-2}),\quad(\nabla_{\tilde{\rho}}\pi)(\tilde{\rho},\tilde{\rho})=o(|\tilde{x}|^{-3}),\quad k(\cdot,\tilde{\rho})=o(|\tilde{x}|^{-2}),$
(84) $|k^{\text{even}}|(\tilde{x})+|\tilde{x}||(\nabla k)^{\text{odd}}|(\tilde{x})=\mathcal{O}(|\tilde{x}|^{-3}).$
@@
(88) $\begin{split}|\xi|^{-1}\xi^{i}(\partial_{i}G_{r,3})(\xi)=&-r^{2}\int_{S_{r}(q)}2W_{2}\alpha d\mu_{\delta}-r^{2}\int_{S_{r}(q)}2\mathcal{D}W_{2}\left(r\delta(\frac{\xi}{|\xi|},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}\\ &-\frac{r^{2}}{2}\bar{D}^{2}\left(\int_{S_{r}(\xi r)}2W_{2}\alpha d\mu_{\delta}\right)_{|\xi=\tilde{\xi}}(r\xi-q,r\xi-q).\\ =&\mathcal{O}(r^{-1})+o(r^{-1}).\end{split}$
@@
Now proceeding as before and using that $\frac{r^{3}}{128\pi}\int_{S_{r}(q)}2\mathcal{D}W_{2}\left(r\delta(\frac{\xi}{|\xi|},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}=o(1)$ we have
(89) $\begin{split}r\delta(\xi(r),&e_{i})=\delta(C_{H},e_{i})+\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\mathrm{Sc}\,\delta(e_{i},\nu)+r\pi(\nu,\nu)\nabla_{\nu}\pi(\nu,\nu)\delta(\nu,e_{i})+2\pi(\nu,\nu)k(\nu,e_{i})\\ &+\pi(\nu,\nu)^{2}\delta(\nu,e_{i})\,d\mu_{\delta}+\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\Big(\delta(e_{i},\nu)\delta(r\xi-q,\nu)\bar{D}_{\nu}\mathrm{Sc}+3\delta(e_{i},\nu)\delta(r\xi-q,\nu)\frac{\mathrm{Sc}}{r}\\ &-\frac{\mathrm{Sc}}{r}\,\delta(e_{i},r\xi-q)\Big)d\mu_{\delta}+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}2\mathcal{D}W_{2}\left(\delta(e_{i},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}+\mathcal{O}(r^{-1})\\ &\,{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}11}=\delta(C_{H}+Z,e_{i})+\mathcal{O}(r^{-1})+o(1)\end{split}$
@@
(92) $|k^{\text{even}}|(\tilde{x})+|\tilde{x}||(\nabla k)^{\text{odd}}|(\tilde{x})=o(|\tilde{x}|^{-3}).$
@@
(93) $\lim_{r\to\infty}\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\mathrm{Sc}\,\nu^{j}+2r\mathcal{D}W_{2}\left(\delta(e_{j},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}=Z^{j}$
@@
Remark 3.12 (Comparison of the Hawking energy).
It follows from the decay of the first variation of the Hawking functional that the Hawking energy of the leaves of the foliation can be estimated by that of coordinate spheres. More precisely, one has
(95) $\mathcal{E}(S_{r}(r\xi))-\mathcal{E}(\Sigma(r))=\mathcal{O}(r^{-2})\quad\text{and}\quad\mathcal{E}(S_{r}(0))-\mathcal{E}(S_{r}(r\xi))=\mathcal{O}(|\xi|),$
so that
(96) $\mathcal{E}(\Sigma(r))=\mathcal{E}(S_{r}(0))+\mathcal{O}(|\xi|).$
Under the assumptions of Section 2, the translation parameter $|\xi|$ is controlled by the decay of the scalar curvature and of the tensor $k$. In particular, under the stronger asymptotic conditions of Section 3, one has $|\xi|=\mathcal{O}(r^{-1})$, and hence
(97) $\mathcal{E}(\Sigma(r))=\mathcal{E}(S_{r}(0))+\mathcal{O}(r^{-1}).$
@@ 4.1. Nonnegativity along the foliation
In [rigidiaz], the following nonnegativity and rigidity result for Hawking surfaces was obtained.
Theorem 4.1 ([rigidiaz, Theorem 3.6]).
@@
As mentioned in the remark following [rigidiaz, Theorem 3.6] one could define $f$ differently,
@@
First, we note that the Hawking energy is positive both along the foliation of Theorem 2.10 and for the on–center family of Theorem 2.14. This follows directly from [diaz2023local, Theorem 2.2] (which treats the foliation) or from a direct application of Theorem 4.1.
@@ 4.2. Monotonicity
Monotonicity lies at the very heart of what one expects from a physically meaningful quasi‐local energy: as a family of surfaces expands outward in an initial data set satisfying the dominant energy condition, its measured energy should not decrease. One of the most celebrated applications of monotonicity properties is in the proof of the Riemannian Penrose inequality by Huisken and Ilmanen [huisken2001inverse]. In their work, they employed the inverse mean curvature flow to show that the Hawking energy is nondecreasing along the flow under suitable conditions, such as nonnegative scalar curvature. This monotonicity, in the time-symmetric initial data setting, allowed them to rigorously relate the ADM mass of an asymptotically flat manifold to the area of its outermost apparent horizon.
Since then, there have been several generalizations of this monotonicity for the Hawking energy to different settings. For instance, Bray, Jauregui, and Mars in [bray2015time, bray2016time] established monotonicity results for the spacetime setting for "time flat" surfaces, while more recently, Hirsch studied monotonicity in a general initial data set setting [hirsch2022hawking].
@@
We proceed as in [willflat, Theorem 4]. We compute the variation of the Hawking energy
@@
In the early development of the conformal method for constructing solutions to the Einstein constraint equations, York studied the asymptotically flat case in [york1979kinematics] and proposed the following model for the extrinsic curvature:
@@
When deriving this model in Chapter 6 of York’s paper [york1979kinematics], the transverse-traceless (TT) part of the tensor $A$, denoted by $A^{*}$, is not explicitly determined during the construction of the model for asymptotically flat manifolds. Instead, York focuses on the decomposition of $A$ into its TT part and a divergence-free component derived from the vector potential $W$:
@@ 4.3. Rigidity on the foliation
In [rigidiaz], the following rigidity result for the Wilmore foliation was obtained.
Theorem 4.11 ([rigidiaz, Theorem 2.16]).
@@
Denoting by $\Omega$ interior of $\Sigma$, we can directly apply Theorem 4.1 obtaining that $\Omega$ is isometric to a spacelike hypersurface in Minkowski spacetime, also we have that $\Sigma$ is a round sphere with $|k|_{\Sigma}=\lambda=|\mathring{B}|=\mathrm{Sc}_{|\Sigma}=\mathrm{Ric}_{|\Sigma}(\nu,\nu)=0$ and $H=\frac{2}{r^{2}}$ is constant, where $r$ is the area radius of $\Sigma$ (for more details see [rigidiaz, Theorem 3.6]). Then what it remains to see is that $M\setminus\Omega$ is also isometric to a hypersurface in Minkowski space.
@@ Appendix A Density of Harmonic asymptotics
Corvino and Schoen showed in [CorvinoSchoen, Theorem 1] that initial data sets with harmonic asymptotics are dense in the set of solutions of the set of asymptotically flat vacuum solutions of the constrained equations. This result was later generalized to solutions of the Einstein constraint equations under the dominant energy condition by Eichmair, Huang and Lee in [eichmair2015spacetime, Theorem 18], this result can be stated assuming higher regularity like in Definition A.3 (the proof of the result is basically unchanged).
@@
Note that in particular, we have also a density result in the Hölder norm by the following result which can be found in [lee2021geometric, Theorem A.25]:
@@ Acknowledgements.
The Author would like to thank Jan Metzger, Marc Mars, Rodrigo Avalos and Thomas Koerber for the helpful comments and discussions about this work. The author also thanks the anonymous referee for helpful and constructive comments that improved this paper. This research received partial support from the DFG as part of the SPP 2026 Geometry at infinity, project ME 3816/3-1.